# Asymptotic expansions for Legendre functions via differential equations having coalescing turning points

### v2

@@ 4 Expansions valid at the turning points
This surrounds, but does not contain, the interval $[z_{1},z_{2}]$. The significance of $D$ is that any numerically satisfactory pair of LG expansions is asymptotically valid in this domain. We also need to utilize LG expansions for parabolic cylinder functions, and as such we also need to exclude the points where they are not all valid, namely those corresponding to the $\zeta$ interval $[-\hat{\alpha},\hat{\alpha}]$. Now, we know that $\zeta(z_{1})=-\alpha$ and $\zeta(z_{2})=\alpha$. This leads to the following.
@@
Note that $J$ would be the interval $[\hat{z}_{1},\hat{z}_{2}]$ if $\hat{\alpha}$ were real. We also observe that, for large $u$, the point sets $[z_{1},z_{2}]$ and $J$ are close to each other, since from Eqs. 37 and 84 $\hat{\alpha}=\alpha+\mathcal{O}(u^{-2})$ as $u\to\infty$.
For $z\in D\setminus I(\delta)$ we can then use both the LG expansions of the general equation given in Section 2, as well as those for parabolic cylinder functions given in Section 3. Thus, we can use the appropriate LG expansions from Eqs. 18, 19, 20, 28, 29, 59, 60, 61, 65, 66, 67, 68, 74, 75, 76, 77, 78, 79, and 80 in Eqs. 86 and 87, and as a result we find that the same expansions are obtained regardless of which pair is chosen, namely
(89) $\mathcal{A}(u,a,z)\sim\frac{(2u)^{1/4}}{2}\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\left[K_{n,2}(u,a)\exp\left\{u(\hat{\xi}-\xi)+\sum_{s=1}^{\infty}(-1)^{s}\frac{\tilde{\mathcal{E}}_{s}(a,z)}{u^{s}}\right\}\right.\\ \left.-K_{n,1}(u,a)\exp\left\{-u(\hat{\xi}-\xi)+\sum_{s=1}^{\infty}\frac{\tilde{\mathcal{E}}_{s}(a,z)}{u^{s}}\right\}\right],$
@@
(90) $\mathcal{B}(u,a,z)\sim-\frac{1}{(2u)^{3/4}}\frac{1}{\left\{f(a,z)\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\\ \times\left[K_{n,2}(u,a)\exp\left\{u(\hat{\xi}-\xi)+\sum_{s=1}^{\infty}(-1)^{s}\frac{\mathcal{E}_{s}(a,z)}{u^{s}}\right\}\right.\\ \left.+K_{n,1}(u,a)\exp\left\{-u(\hat{\xi}-\xi)+\sum_{s=1}^{\infty}\frac{\mathcal{E}_{s}(a,z)}{u^{s}}\right\}\right],$
as $u\to\infty$, both of which hold uniformly for $z\in D\setminus I(\delta)$. Here
(91) $K_{n,1}(u,a)=i\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{-\frac{1}{4}u\hat{\alpha}^{2}}k_{n,1}(u,a),$
@@
for $\alpha\in[0,\alpha_{1}]$ and bounded $\zeta$. It can also be confirmed that $\hat{\xi}-\xi=u^{-2}\mathcal{O}\{\ln(\zeta)\}+\mathcal{O}(u^{-1})$ as $\zeta\to\infty$. Therefore, the terms ±$\pm u(\hat{\xi}-\xi)$ appearing in the exponents in Eqs. 89 and 90 are bounded for fixed $\zeta$ as $u\to\infty$, and contribute at worst (on exponentiation) to a slowly growing algebraic term as $\zeta\to\infty$ with $u$ large. This confirms that $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ are slowly varying as $u\to\infty$ and $z\in D\setminus I(\delta)$.
@@
The contour $\mathcal{C}$ is a suitably chosen positively orientated simple loop lying in $D$ that encloses $z$ and the set $I(\delta)$. In the integrands we can then insert the expansions Eqs. 89 and 90, since they are valid on the contour. This enables an accurate evaluation of $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ for points inside the contour, particularly those in $I(\delta)$.
Note that Eqs. 81, 82, 83, and 84, being exact, can be differentiated with respect to $z$ to produce expansions for the derivatives of the solutions. In doing so, one uses the chain rule in conjunction with (5), as well as 13). For example, from (81)
(97) $\frac{\partial w_{n,1}(u,a,z)}{\partial z}=e^{-\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\Biggl{\{}U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)\tilde{\mathcal{A}}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\tilde{\mathcal{B}}(u,a,z)\right\},$
@@
Suppose that the odd coefficients $\hat{E}_{2s+1}(a,z)$ are analytic in $D$, for example, being rational functions of $z$ and $\sigma(z):=\{(z-z_{1})(z-z_{2})\}^{1/2}$. Also suppose for $a_{0}\leq a\leq a_{1}$ that
(102) $\frac{k_{n,1}(u,a)}{k_{n,2}(u,a)}\sim\sum_{s=0}^{\infty}\frac{\kappa_{s}(a)}{u^{s}}\quad(u\to\infty),$
for certain constants $\kappa_{s}(a)$ with $\kappa_{0}(a)\neq 0$ (cf. Lemma 2.3). Next, assume for some fixed $z_{0}\in\mathbb{C}$, positive $\rho_{0}$, and sufficiently small $\delta>0$, that the disk $D_{0}=\{z:\,|z-z_{0}|<\rho_{0}\}$ contains $I(\delta)$ and lies inside $D\cup I(\delta)$. Then as $u\to\infty$, $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ possess the asymptotic expansions
(103) $\mathcal{A}(u,a,z)\sim\Lambda_{1}(u,a)\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\left\{1+\sum_{s=1}^{\infty}\frac{\mathrm{A}_{s}(a,z)}{u^{2s}}\right\},$
@@
(104) $\mathcal{B}(u,a,z)\sim\Lambda_{2}(u,a)\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\sum_{s=0}^{\infty}\frac{\mathrm{B}_{s}(a,z)}{u^{2s+2}},$
@@
for certain assignable constants $\Lambda_{j}(u,a)$ ($j=1,2$), uniformly for $a_{0}\leq a\leq a_{1}$ and $z\in D\cup I(\delta)$. The coefficients $\mathrm{A}_{s+1}(a,z)$ and $\mathrm{B}_{s}(a,z)$ ($s=0,1,2,\ldots$) can be evaluated by expanding Eqs. 89 and 90 into the forms Eqs. 103 and 104, and are analytic in $D\cup I(\delta)$.
Proof 4.4.
We know that the even coefficients $\hat{E}_{2s}(a,z)$ are always meromorphic at $z_{1}$ and $z_{2}$, and therefore are analytic in $D$. Now, by hypothesis, the odd coefficients $\hat{E}_{2s+1}(a,z)$ are also analytic in $D$. Hence from Theorems 2.1 and 84 it follows that $\alpha=\hat{\alpha}+\mathcal{O}(u^{-n})$ for any positive integer $n$, and therefore, similarly to the derivation of (95), $\xi-\hat{\xi}=\mathcal{O}(u^{-n})$ for $\zeta$ bounded, again for any positive integer $n$.
@@
Consider now an annulus $R=\{z:\,\rho_{0}<|z-z_{0}|<\rho_{1}\}$ that lies within $D$, with $\rho_{0}$ defined by the hypothesis of the theorem. From this definition $I(\delta)$ lies inside $D_{0}$ for sufficiently small positive $\delta$, and therefore does not lie inside $R$. It follows from Eqs. 93 and 94 that ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are analytic in $R$ (see also the comments after Eqs. 58 and 73.
Now for $z\in R$ we have by hypothesis, as noted above, $\xi-\hat{\xi}=\mathcal{O}(u^{-n})$ for any positive integer $n$, and consequently one can show, using Eqs. 40 and 102 and [3, Eq. 5.11.1], that Eqs. 89 and 90 can indeed be expanded into formal asymptotic series of the forms Eqs. 103 and 104. Furthermore, each $\mathrm{A}_{s}(a,z)$ and $\mathrm{B}_{s}(a,z)$ must be analytic in the annulus $R$, since, as we have shown, the same is true of all the coefficients appearing in Eqs. 89 and 90.
@@
To extend this to $D_{0}$, and in particular $I(\delta)$, we appeal to Theorem 4.2. From this it follows under the above assumptions that the coefficients $\mathrm{A}_{s}(a,z)$ and $\mathrm{B}_{s}(a,z)$ are actually analytic in $D_{1}=\{z:\,|z-z_{0}|<\rho_{1}\}$, and in particular at both turning points (whether separated or not), and moreover the asymptotic expansions Eqs. 103 and 104 hold in this disk, and by extension to all points in $D\cup I(\delta)$.
@@
Let us summarize the main results of this section. Consider the differential equation Eq. 1 with the assumptions on $f(a,z)$ and $g(z)$ stated in the paragraph following that equation, and in particular, for $a_{0}\leq a\leq a_{1}$ $f(a,z)$ has simple zeros at $z=z_{1}$ and $z=z_{2}$, which coalesce as $a\to a_{0}$. Define $\xi$, $\zeta$, and $\alpha$ by Eqs. 2, 8, and 11. Let coefficients $\hat{E}_{s}(a,z)$ be given by Eqs. 21, 22, and 23. For each positive integer $n$ let $w_{n,j}(u,a,z)$ ($j=1,2,3,4$) be solutions of Eq. 1 given by Eqs. 18, 19, 20, 28, and 29, which are recessive at the singularities $z=x_{j}$ ($j=1,2,3,4$) of Eq. 1 which correspond to $\zeta=i\infty$, $\zeta=+\infty$, $\zeta=-i\infty$, and $\zeta=-\infty$, respectively. Next $k_{n,j}(u,a)$ ($j=1,2,3$) and $k_{n,4}^{\pm}(u,a)$ are constants, continuous for $a_{0}\leq a\leq a_{1}$, chosen such that the connection formulas Eqs. 30 and 31 hold for some $\lambda$. Define $\hat{\xi}$ and $\hat{\beta}$ by Eqs. 49, 52, and 53, where $\hat{\alpha}$ is given by Eq. 84, and let coefficients $e_{s}(\hat{\alpha},\hat{\beta})$ and $\tilde{e}_{s}(\hat{\alpha},\hat{\beta})$ be given by Eqs. 56, 57, 72, and 73, with subsequent ones recursively given by Eq. 58. Let $D$ be a domain that surrounds, but does not contain, the set $I(\delta)$ that contains the turning points (defined by Definition 4.1), and is given by Eq. 88 along with certain LG regions of validity described at the beginning of Section 2.
@@
Then $w_{n,j}(u,a,z)$ can be expressed in the forms Eqs. 81, 82, 83, and 85, where, for $a_{0}\leq a\leq a_{1}$ and $z\in D$, $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ possess the asymptotic expansions Eqs. 89 and 90. In these $K_{n,1}(u,a)$, $K_{n,2}(u,a)$, ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are given by Eqs. 91, 92, 93, and 94.
@@
For points in $I(\delta)$, which contains the two turning points, the coefficient functions $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ can be evaluated via the Cauchy integrals Eq. 96, with the expansions Eqs. 89 and 90 used on the contour, which is a simple positively orientated loop in $D$ which surrounds $z$ and both turning points.
@@
The expansions Eqs. 89 and 90 can be expanded in the forms Eqs. 103 and 104, under conditions given in Theorem 4.3. If applicable, this gives an alternative way to asymptotically evaluate $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ in $I(\delta)$.

### v3

@@ 4 Expansions valid at the turning points
We assume here that $\{x_{1},x_{2}\}\in Z_{1}\cap Z_{2}$, which means that every point in this intersection lies on a progressive path linking $x_{1}$ to $x_{2}$. Similarly for the other three intersection sets.
Now $D$ surrounds, but does not contain, the interval $[z_{1},z_{2}]$. The significance of $D$ is that any numerically satisfactory pair of LG expansions is asymptotically valid in this domain. We also need to utilize LG expansions for parabolic cylinder functions, and as such we also need to exclude the points where they are not all valid, namely those corresponding to the $\zeta$ interval $[-\hat{\alpha},\hat{\alpha}]$. Now, we know that $\zeta(z_{1})=-\alpha$ and $\zeta(z_{2})=\alpha$. This leads to the following.
@@
Note that $J$ would be the interval $[\hat{z}_{1},\hat{z}_{2}]$ if $\hat{\alpha}$ were real. We also observe that, for large $u$, the point sets $[z_{1},z_{2}]$ and $J$ are close to each other, since from Eqs. 37 and 89 $\hat{\alpha}=\alpha+\mathcal{O}(u^{-2})$ as $u\to\infty$.
For $z\in D\setminus I(\delta)$ we can then use both the LG expansions of the general equation given in Section 2, as well as those for parabolic cylinder functions given in Section 3. Thus, we can use the appropriate LG expansions from Eqs. 2, 18, 19, 20, 28, 29, 53, 63, 64, 65, 70, 71, 72, 73, 79, 80, 81, 82, 83, 84, and 85 in Eqs. 91 and 92, and as a result we find that the same expansions are obtained regardless of which pair is chosen, namely
(94) $\mathcal{A}(u,a,z)=\frac{(2u)^{1/4}}{2}\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\\ \times\left[K_{n,2}(u,a)\exp\left\{u(\hat{\xi}-\xi)+\sum_{s=1}^{n-1}(-1)^{s}\frac{\tilde{\mathcal{E}}_{s}(a,z)}{u^{s}}\right\}\right.\\ \left.-iK_{n,1}(u,a)\exp\left\{-u(\hat{\xi}-\xi)+\sum_{s=1}^{n-1}\frac{\tilde{\mathcal{E}}_{s}(a,z)}{u^{s}}\right\}\right]\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\},$
@@
(95) $\mathcal{B}(u,a,z)=-\frac{1}{(2u)^{3/4}}\frac{1}{\left\{f(a,z)\left(\zeta^{2}-\hat{\alpha}^{2}\right)\right\}^{1/4}}\\ \times\left[K_{n,2}(u,a)\exp\left\{u(\hat{\xi}-\xi)+\sum_{s=1}^{n-1}(-1)^{s}\frac{\mathcal{E}_{s}(a,z)}{u^{s}}\right\}\right.\\ \left.+iK_{n,1}(u,a)\exp\left\{-u(\hat{\xi}-\xi)+\sum_{s=1}^{n-1}\frac{\mathcal{E}_{s}(a,z)}{u^{s}}\right\}\right]\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\},$
as $u\to\infty$, uniformly for $z\in D\setminus I(\delta)$, where
(96) $K_{n,1}(u,a)=\left(\frac{u\hat{\alpha}^{2}}{2e}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}k_{n,1}(u,a),$
@@
for $\alpha\in[0,\alpha_{1}]$ and bounded $\zeta$. It can also be confirmed that $\hat{\xi}-\xi=u^{-2}\mathcal{O}\{\ln(\zeta)\}+\mathcal{O}(u^{-1})$ as $\zeta\to\infty$. Therefore, the terms ±$\pm u(\hat{\xi}-\xi)$ appearing in the exponents in Eqs. 94 and 95 are bounded for fixed $\zeta$ as $u\to\infty$, and contribute at worst (on exponentiation) to a slowly growing algebraic term as $\zeta\to\infty$ with $u$ large. This confirms that $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ are slowly-varying as $u\to\infty$ and $z\in D\setminus I(\delta)$.
@@
The contour $\mathcal{C}$ is a suitably chosen positively orientated simple loop lying in $D$ that encloses $z$ and the set $I(\delta)$. In the integrands we can then insert the expansions Eqs. 94 and 95, since they are valid on the contour. This enables an accurate evaluation of $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ for points inside the contour, particularly those in $I(\delta)$.
Note that Eqs. 86, 87, 88, and 89, being exact, can be differentiated with respect to $z$ to produce expansions for the derivatives of the solutions. In doing so, one uses the chain rule in conjunction with Eq. 5, as well as Eq. 13. For example, from Eq. 86
(102) $\frac{\partial w_{1}(u,a,z)}{\partial z}=e^{-\frac{1}{4}(u\hat{\alpha}^{2}-1)\pi i}\Biggl{\{}U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)\tilde{\mathcal{A}}(u,a,z)\Biggr{.}\\ \left.+\frac{\partial U\left(\tfrac{1}{2}u\hat{\alpha}^{2},-i\sqrt{2u}\,\zeta\right)}{\partial\zeta}\tilde{\mathcal{B}}(u,a,z)\right\},$
@@
For arbitrary positive integer $n$, and $a_{0}\leq a\leq a_{1}$, suppose that
(107) $\frac{K_{n,1}(u,a)}{K_{n,2}(u,a)}=l_{0}(a)\exp\left\{\sum_{s=1}^{n-1}\frac{l_{s}(a)}{u^{s}}\right\}\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty),$
for certain constants $l_{2s+1}(a)$ with $l_{0}(a)\neq 0$, where $K_{n,1}(u,a)$ and $K_{n,2}(u,a)$ are defined by Eqs. 96 and 97. Also, assume for some fixed $z_{0}\in\mathbb{C}$, positive $\rho_{0}$, and sufficiently small $\delta>0$, that the disk $D_{0}=\{z:\,|z-z_{0}|<\rho_{0}\}$ contains $I(\delta)$ and lies inside $D\cup I(\delta)$. Then as $u\to\infty$, $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ possess asymptotic expansions of the forms
(108) $\mathcal{A}(u,a,z)\sim\Lambda(u,a)\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\left\{1+\sum_{s=1}^{\infty}\frac{\mathrm{A}_{s}(a,z)}{u^{s}}\right\},$
@@
(109) $\mathcal{B}(u,a,z)\sim\frac{\Lambda(u,a)}{u^{2}}\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\sum_{s=0}^{\infty}\frac{\mathrm{B}_{s}(a,z)}{u^{s}},$
@@
for a certain assignable constant $\Lambda(u,a)$, uniformly for $z\in D\cup I(\delta)$ and $a_{0}\leq a\leq a_{1}$. In addition, the coefficients $\mathrm{A}_{s+1}(a,z)$ and $\mathrm{B}_{s}(a,z)$ ($s=0,1,2,\ldots$) are analytic in $D\cup I(\delta)$.
Remark 4.4.
From Eq. 10, Eq. 37, Eq. 89, Eq. 95, Eq. 96, Lemma 2.3, and [3, Eq. 5.11.1], it follows that
(110) $\frac{K_{n,1}(u,a)}{K_{n,2}(u,a)}=i\exp\left\{\sum_{s=0}^{n-1}\frac{l_{2s+1}(a)}{u^{2s+1}}\right\}\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty),$
under the weaker condition $0<a_{0}+\delta\leq a\leq a_{1}$.
Proof 4.5.
Firstly, from Eqs. 18, 20, 25, 26, and 27,
(111) $f^{1/4}(a,z)e^{u\xi}w_{2}(u,a,z)\to\gamma_{2}(u,a)\quad(z\to x_{2}),$
where $\gamma_{2}(u,a)$ is given by
(112) $\gamma_{2}(u,a)=k_{n,2}(u,a)\exp\left\{\sum\limits_{s=1}^{n-1}{(-1)^{s}\frac{\hat{E}_{s}(a,x_{2})}{u^{s}}}\right\}\left\{1+\eta_{n,2}(u,a,x_{2})\right\},$
for arbitrary positive integer $n$; observe from Eq. 111 that $\gamma_{2}(u,a)$ is independent of $n$. Hence from Eq. 97
(113) $K_{n,2}(u,a)=\frac{1}{\sqrt{2\pi}}\left(\frac{2e}{u\hat{\alpha}^{2}}\right)^{\frac{1}{4}u\hat{\alpha}^{2}}\Gamma\left(\tfrac{1}{2}u\hat{\alpha}^{2}+\tfrac{1}{2}\right)\gamma_{2}(u,a)\\ \times\exp\left\{\sum\limits_{s=1}^{n-1}{(-1)^{s-1}\frac{\hat{E}_{s}(a,x_{2})}{u^{s}}}\right\}\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}\quad(u\to\infty).$
Now rewrite Eq. 94 in the form
(114) $\mathcal{A}(u,a,z)=\frac{(2u)^{1/4}}{2}K_{n,2}(u,a)\left(\frac{\zeta^{2}-\hat{\alpha}^{2}}{f(a,z)}\right)^{1/4}\\ \times\left[\exp\left\{u(\hat{\xi}-\xi)+\sum_{s=1}^{n-1}(-1)^{s}\frac{\tilde{\mathcal{E}}_{s}(a,z)}{u^{s}}\right\}\right.\\ \left.-i\frac{K_{n,1}(u,a)}{K_{n,2}(u,a)}\exp\left\{-u(\hat{\xi}-\xi)+\sum_{s=1}^{n-1}\frac{\tilde{\mathcal{E}}_{s}(a,z)}{u^{s}}\right\}\right]\left\{1+\mathcal{O}\left(\frac{1}{u^{n}}\right)\right\}.$
Then, for $z\in D\setminus I(\delta)$ and $a_{0}\leq a\leq a_{1}$, one can show, using Eqs. 2, 53, 37, 41, 107, and 113, that Eq. 94 can be expanded into a formal asymptotic series of the form Eq. 108. The expansion Eq. 109 can be obtained in a similar way.
Next, it follows from Eqs. 2, 8, 53, 98, and 99 (and see also the comments after Eqs. 62 and 78) that $\hat{\xi}$, $\xi$, ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are analytic in $D$ except the first two (regarded as functions of $z$), and possibly the odd coefficients of the latter two, having points of discontinuity on the real axis to the left of $z=z_{1}$. As such, all are certainly analytic for $z\in D\setminus\{[x_{4},z_{1}]\cup I(\delta)\}$.
@@
Consider now an annulus $R=\{z:\,\rho_{0}<|z-z_{0}|<\rho_{1}\}\subset D$, with $\rho_{0}$ defined by the hypothesis of the theorem. The definition of $\rho_{0}$ implies that $I(\delta)$ lies inside $D_{0}$ for sufficiently small positive $\delta$, and therefore does not lie within $R$. Thus, from above, $\hat{\xi}$, $\xi$, ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are analytic in $R\setminus[-\rho_{1},-\rho_{0}]$, and are bounded in $R$.
Furthermore, from the re-expansion of Eq. 114 into Eq. 108, and similarly from the construction of Eq. 109, which are both valid in $R$, we infer that each $\mathrm{A}_{s}(a,z)$ and $\mathrm{B}_{s}(a,z)$ is a polynomial in terms of $\hat{\xi}$, $\xi$, ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$. Thus they too must be analytic in the cut annulus $R\setminus[-\rho_{1},-\rho_{0}]$, and bounded in $R$. However, $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ are analytic in $R$, and in particular continuous at $[-\rho_{1},-\rho_{0}]$, and so from Eqs. 108 and 109 we conclude that $\mathrm{A}_{s}(a,z)$ and $\mathrm{B}_{s}(a,z)$ must also be continuous at this interval. Accordingly, these coefficients are analytic at the cut, and hence throughout $R$.
@@
To further extend this to $D_{1}=\{z:\,|z-z_{0}|<\rho_{1}\}$, and in particular to $I(\delta)$, we appeal to Theorem 4.2. From this it follows under the above assumptions that the coefficients $\mathrm{A}_{s}(a,z)$ and $\mathrm{B}_{s}(a,z)$ are actually analytic in $D_{1}$, which contains both turning points (whether separated or not). Moreover, the asymptotic expansions Eqs. 108 and 109 hold in this disk, and by extension to all points in $D\cup I(\delta)$.
@@
Let us summarize the main results of this section. Consider the differential equation Eq. 1 with the assumptions on $f(a,z)$ and $g(z)$ stated in the paragraph following that equation, and in particular, for $a_{0}\leq a\leq a_{1}$ $f(a,z)$ has simple zeros at $z=z_{1}$ and $z=z_{2}$, which coalesce as $a\to a_{0}$. Define $\xi$, $\zeta$, and $\alpha$ by Eqs. 2, 8, and 11. Let coefficients $\hat{E}_{s}(a,z)$ be given by Eqs. 21, 22, and 23. For each positive integer $n$ let $w_{j}(u,a,z)$ ($j=1,2,3,4$) be solutions of Eq. 1 given by Eqs. 18, 19, 20, 28, and 29, which are recessive at the singularities $z=x_{j}$ ($j=1,2,3,4$) of Eq. 1 which correspond to $\zeta=i\infty$, $\zeta=+\infty$, $\zeta=-i\infty$, and $\zeta=-\infty$, respectively. Next, $k_{n,j}(u,a)$ ($j=1,2,3$) and $k_{n,4}^{\pm}(u,a)$ are constants, continuous for $a_{0}\leq a\leq a_{1}$, chosen such that the connection formulas Eqs. 30 and 31 hold for some $\lambda$. Define $\hat{\xi}$ and $\hat{\beta}$ by Eqs. 53, 56, and 57, where $\hat{\alpha}$ is given by Eq. 89, and let coefficients $e_{s}(\hat{\alpha},\hat{\beta})$ and $\tilde{e}_{s}(\hat{\alpha},\hat{\beta})$ be given by Eqs. 60, 61, 77, and 78, with subsequent ones given recursively by Eq. 62. Let $D$ be a domain that surrounds, but does not contain, the set $I(\delta)$ that contains the turning points (defined by Definition 4.1), and is given by Eq. 93 along with certain LG regions of validity described at the beginning of Section 2.
@@
Then $w_{j}(u,a,z)$ can be expressed in the forms Eqs. 86, 87, 88, and 90, where, for $a_{0}\leq a\leq a_{1}$ and $z\in D$, $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ possess the asymptotic expansions Eqs. 94 and 95. In these $K_{n,1}(u,a)$, $K_{n,2}(u,a)$, ~$\tilde{\mathcal{E}}_{s}(a,z)$ and $\mathcal{E}_{s}(a,z)$ are given by Eqs. 96, 97, 98, and 99.
@@
For points in $I(\delta)$, which contains the two turning points, the coefficient functions $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ can be evaluated via the Cauchy integrals Eq. 101, with the expansions Eqs. 94 and 95 used on the contour, which is a simple positively orientated loop in $D$ which surrounds $z$ and both turning points.
@@
The expansions Eqs. 94 and 95 can be expanded in the forms Eqs. 108 and 109, under conditions given in Theorem 4.3. If applicable, this gives an alternative way to asymptotically evaluate $\mathcal{A}(u,a,z)$ and $\mathcal{B}(u,a,z)$ in $I(\delta)$.